# Transferable Ensemble Black-box Jailbreak Attacks on Large Language Models

### v1

@@ 1 Introduction
Jailbreaking large language models (LLMs) is being intensively studied to evaluate the safety of LLMs [1, 2]. Existing jailbreaking methods, as adversarial attacks targeting small-scale learning models, can be categorized into white-box and black-box approaches. In white-box methods, gradient-based optimization techniques are employed to identify adversarial suffixes [3]. For black-box methods, various optimization strategies, such as genetic algorithms, are used to refine jailbreak prompt templates through rephrasing, word replacement, and other modifications [4, 5]. Recent research [6] has demonstrated that LLMs can function as powerful optimizers when provided with sufficient contextual information. Indeed, both the persuasive adversarial prompt (PAP) [7] and Tree of Attacks with Pruning (TAP) [8] methods leverage multiple LLMs to optimize malicious instructions. In this line of research, given a target aligned LLM to jailbreak, one LLM operates as an attacker, optimizing a malicious instruction to circumvent the unknown safety mechanisms of the target LLM. A second LLM serves as a judge, observing the interaction history between the attacker and the target LLM, evaluating the attack’s effectiveness, and providing feedback to the attacker LLM. The PAP method [7] has been officially adopted as the baseline attack strategy in the Competition for LLM and Agent Safety 2024 (CLAS 2024) [^1] https://www.llmagentsafetycomp24.com/.

### v2

@@ 1 Introduction
Jailbreaking large language models (LLMs) is being intensively studied to evaluate the safety of LLMs [1, 2]. Existing jailbreaking methods, as adversarial attacks targeting small-scale learning models, can be categorized into white-box and black-box approaches. In white-box methods, gradient-based optimization techniques are employed to identify adversarial suffixes [3]. For black-box methods, various optimization strategies, such as genetic algorithms, are used to refine jailbreak prompt templates through rephrasing, word replacement, and other modifications [4, 5]. Recent research [6] has demonstrated that LLMs can function as powerful optimizers when provided with sufficient contextual information. Indeed, both the persuasive adversarial prompt (PAP) [7] and Tree of Attacks with Pruning (TAP) [8] methods leverage multiple LLMs to optimize malicious instructions. For the PAP method [7], two LLMs are utilized for the whole jailbreak optimization and for the TAP method [8], three LLMs are utilized. The PAP method [7] has been officially adopted as the baseline attack strategy in the Competition for LLM and Agent Safety 2024 (CLAS 2024) [^1] https://www.llmagentsafetycomp24.com/.